OpenStudy (anonymous):

What are 5 names for the number 8 that include the fraction 1/3 and multiplication ?

OpenStudy (anonymous):

names??

OpenStudy (anonymous):

As in equations

OpenStudy (anonymous):

but they have to include the following : the fraction 1/3 and multiplication

OpenStudy (anonymous):

do you mean like 8=(x/3) where x=24 something like that??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

well it should be quite easy to make up equations like this, so if they have to = 8 then you could have: 8=(x/3) + (2y) where x could = 6 and y would the have to = 3 there are load of possibilities

OpenStudy (anonymous):

soz y=3 i cant type

OpenStudy (anonymous):

ok thanx